OpenStudy (anonymous):

Find the factors of 2a2 + 9a − 18. (2a − 3)(a + 6) (2a + 3)(a − 6) (2a − 9)(a + 2) (2a − 2)(a + 9)

OpenStudy (swissgirl):

hmm how abt we dont do the guessing game and just solve it

OpenStudy (swissgirl):

A=2 B=9 and C=-18 So we need to find 2 numbers that when you multiply it equals A*C=2*18=36 and when you add these 2 numbers you get B=9

OpenStudy (anonymous):

Yup. Pick an answer choice. Multiply it out and see if you get \(2a^2 +9a -18\)

OpenStudy (anonymous):

i got (2a + 3)(a − 6)

OpenStudy (swissgirl):

Close enough. You gotta sharpen your guessing skills :P

OpenStudy (anonymous):

Do you know how to multiply that out? Remember FOIL

OpenStudy (mertsj):

@brianasosaa You will find many people in this world that you may think of as mean. If you want them to help you, you might have to overlook it.

OpenStudy (swissgirl):

Are you referring to me ? @Mertsj :P

OpenStudy (radar):

Close, but no cigar! Check out the signs.

OpenStudy (radar):

It would of been correct if the expression had of been: 2a^2 - 9a -18 but alas that was not the case.